Given the data below test the hypothesis that the mean of three populations are equal.let alpha=0.05



Sample-1 sample-2 sample-3



40 70 45



50 65 38



60 66 60



65 50 42

1
Expert's answer
2022-01-04T18:09:54-0500

Solution:

To test the mean of three population we use Analysis of Variance (ANOVA).It is used for testing the equality of three or more population.

μ1 - Mean of population 1

μ2 - Mean of population 2

μ3 - Mean of population 3

Step 1: Formulate the null and alternative hypothesis.

The null Hypothesis (H0) = μ1 =μ2 =μ3

Alternative Hypothesis(H1) = at least one of the means is significantly different from the others.



Σ XA = 215, Σ XB = 251,   Σ XC = 185



Σ XA2 = 11925 , Σ XB2 = 15981, Σ XC2 = 8833

Let k be the number of different samples = 3

n = Total number of samples = n1+n2+n3 =4+4+4 = 12

Where n1-Number of samples in population A

          n2 -Number of samples in population B

          n3 - Number of samples in population C

Data table



sum of squares between samples

(∑Ti2/ni)= (215)2/4+(251)2/4 +(185)2/4 = 11556.25 + 15750.25 + 8556.25 = 35862.75


SSB = (∑Ti2/ni) - (∑x)2/n =    35862.75 - 423801/12 = 35862.75 - 35316.75 = 546

sum of squares within samples

SSW = ∑x2-(∑Ti2/ni) = 36739 - 35862.75 = 876.25

Total sum of squares

SST=SSB+SSW

      = 546 + 876.25

SST = 1422.25

variance between samples

MSB=SSB/(k-1)

MSB = 546/(3-1) = 546/2 = 273

variance within samples

MSW = SSW/(n-k)

MSW  = 876.25 /9 = 97.36

test statistic F for one way ANOVA test

F=MSB/MSW

F = 273/97.36

F =2.80

The degree of freedom between samples

k-1=2

The degree of freedom within samples

n-k = 12-3=9

ANOVA table



F(k -1,n-k) = F(2,9) at 0.05 level of significance = 4.256

Calculated F value = 2.80

Table value of F = 4.256

Caculated value < Table value

So, H0 is accepted, Hence there is no significant difference between mean of three population.